OpenStudy (anonymous):

i messed up the problem. its, 3(x-7) - x = 2x - 21 the choices are; 7. -2. infinitely many solutions. no solution.

OpenStudy (anonymous):

@nfcfox

OpenStudy (anonymous):

is it still 7?

OpenStudy (nfcfox):

@allismith2953 Yes because if we plug in 7 on both sides, we get 3(7-7)-7=2(7)-21 Which is 3(0)-7=14-21 which is -7=-7 therefore it will be 7. It's not 2 because they won't equal each other, the answer does exist and it is not infinite answers.